# Various Lengths, Constant Speed: Efficient Language Modeling with Lightning Attention

### v1

@@ 3.2 Linear Attention with Tiling
$\textstyle\mathbf{X}=\left[\begin{matrix}\mathbf{X}_{1}\\ \mathbf{X}_{2}\end{matrix}\right],\mathbf{X}_{1}\in\mathbb{R}^{m\times d},\mathbf{X}_{2}\in\mathbb{R}^{(n-m)\times d},\mathbf{X}\in\{\mathbf{Q},\mathbf{K},\mathbf{V}\}.$
@@ 5.1 Lightning Attention Evaluation
Table 2: Performance Comparison on Commonsense Reasoning and Aggregated Benchmarks. For a fair comparison, we report competing methods’ results reproduced by us using their released models. Official results are denoted in italics. Note that TNL-15B only consumes 0.8TB tokens. PS: parameter size (billion). T: tokens (billion). HS: HellaSwag. WG: WinoGrande.
@@ 5.2 TNL Evaluation
Benchmark Results In order to validate the effectiveness of TNL, we pretraining 385M, 1B, 7B, and 15B models on self-collected datasets, the details of the data are in the Appendix D, and tested on Commonsense Reasoning Task, MMLU(Hendrycks et al., 2021) and C-Eval(Huang et al., 2023). For comparison, we selected several open-source models as competitors, including Transformer-based models such as OPT (Zhang et al., 2022), Pythia (Biderman et al., 2023), BLOOM (Workshop et al., 2023), GPT-Neo (Black et al., 2022), Falcon (Almazrouei et al., 2023), LLaMA (Touvron et al., 2023a, b), OpenLLAMA  (Geng & Liu, 2023), Baichuan (Baichuan, 2023), ChatGLM (Zeng et al., 2022; Du et al., 2022), and non-Transformer model RWKV (Peng et al., 2023a). It can be observed in Table 2 that, compared to these models, TNL remains highly competitive.
@@ 5.3 TNL Ablation
GLA Activation Functions: in our study on the GLA (Gated Linear Attention) mechanism, we evaluated activation functions, finding Swish and 1+elu to perform similarly, as detailed in Table  7. However, due to NaN issues with 1+elu in our 7B model, we opted for Swish.

### v2

@@ 3.2 Linear Attention with Tiling
$\begin{gathered}\textstyle\mathbf{X}=\left[\begin{matrix}\mathbf{X}_{1}\\ \mathbf{X}_{2}\end{matrix}\right],\mathbf{X}_{1}\in\mathbb{R}^{m\times d},\mathbf{X}_{2}\in\mathbb{R}^{(n-m)\times d},\\ \mathbf{X}\in\{\mathbf{Q},\mathbf{K},\mathbf{V}\}.\end{gathered}$
@@ 5.1 Lightning Attention Evaluation
Table 2: Performance Comparison on Commonsense Reasoning and Aggregated Benchmarks. For a fair comparison, we report competing methods’ results reproduced by us using their released models. Official results are denoted in italics. PS: parameter size (billion). T: tokens (billion). HS: HellaSwag. WG: WinoGrande.
@@ 5.2 TNL Evaluation
Benchmark Results In order to validate the effectiveness of TNL, we pretraining 385M, 1B, 7B, and 15B models on self-collected datasets, the details of the data are in the Appendix D, and tested on Commonsense Reasoning Task, MMLU(Hendrycks et al., 2021), C-Eval(Huang et al., 2023), and SCROLLS (Shaham et al., 2022). For comparison, we selected several open-source models as competitors, including Transformer-based models such as OPT (Zhang et al., 2022), Pythia (Biderman et al., 2023), BLOOM (Workshop et al., 2023), GPT-Neo (Black et al., 2022), Falcon (Almazrouei et al., 2023), LLaMA (Touvron et al., 2023a, b), OpenLLAMA  (Geng & Liu, 2023), Baichuan (Baichuan, 2023), ChatGLM (Zeng et al., 2022; Du et al., 2022), and non-Transformer model RWKV (Peng et al., 2023a). It can be observed in Table 2 and Table 3 that, compared to these models, TNL remains highly competitive.
@@
  • On SCROLLS (Shaham et al., 2022) benchmark, we assess the large language models trained on a 1 billion parameter and pre-trained using a sequence length of 2048. We present zero-shot performance results for all benchmarks using the LM-Eval-Harness (Gao et al., 2021). For generation tasks within SCROLLS, we employ a greedy search with hyper-parameters top_k set to 5 and top_p set to 1. Our models consistently match or surpass the performance of existing state-of-the-art LLMs in these tasks.
@@ 5.3 TNL Ablation
Table 3: Performance Comparison on SCROLLS (Shaham et al., 2022): A review of models up to 1 billion parameters on 2048 pre-training sequence length. PS: parameter size (billion). T: tokens (billion).
Model PS T GovRep SumScr QMSum Qspr Nrtv QALT CNLI Avg
B B ROUGE-1/2/L ROUGE-1/2/L ROUGE-1/2/L F1 F1 EM EM
OPT 0.35 0.30 2.52/0.53/2.24 7.72/0.68/6.52 8.05/1.79/6.6 13.13 10.13 29.05 9.16 7.55
Pythia 0.40 0.30 4.96/1.19/4.06 2.03/0.2/1.79 7.51/1.43/6.08 15.27 8.24 28.57 15.24 7.43
RWKV 0.43 - 1.63/0.4/1.49 0.94/0.11/0.76 10.19/2.26/8.06 13.16 9.76 26.32 16.49 7.04
TNL 0.39 1.0 3.67/1.16/3.14 8.27/0.82/6.91 13.62/3.29/10.95 14.29 11.69 28.14 17.36 9.48
OPT 1.3 0.3 5.7/2.09/4.41 10.17/0.82/8.29 12.36/3.15/9.85 18.37 13.42 29.15 12.44 10.02
Pythia 1.4 0.3 4.03/1.25/3.33 8.34/0.87/6.97 13.17/3.4/10.92 16.09 11.91 28.72 9.06 9.08
Falcon 1.0 0.35 2.74/0.67/2.37 10.95/1.28/8.66 13.29/3.09/10.58 16.17 12.91 29.19 14.75 9.74
TNL 1.0 1.2 6.81/2.30/5.25 12.28/1.23/9.27 14.60/3.51/11.62 15.02 14.66 28.72 37.32 12.51
@@ B B ROUGE-1/2/L ROUGE-1/2/L ROUGE-1/2/L F1 F1 EM EM
GLA Activation Functions: in our study on the GLA (Gated Linear Attention) mechanism, we evaluated activation functions, finding Swish and 1+elu to perform similarly, as detailed in Table  8. However, due to NaN issues with 1+elu in our 7B model, we opted for Swish.